Swiftsure Ltd has carried forward a deferred tax liability of 170000, arising from differences between
carrying amount and tax bases of the company’s assets. On 30 June 2013, the carrying amounts and tax
bases of the company’s assets were as follows:

Assets Carrying amount Tax base
A $142000 $108000
B $541000 $340000
C $820000 $610000
D $86000 $40000
The current income tax expense for the year ended 30 June 2013 is $8400000 and tax rate is 30%.
Required:
(a) Prepare general journal entries on 30 June 2013 to record company’s income tax expense.
(b) How would your answer differ if the balance of the deferred tax liability carried forward was $120000?
Answer #1

Amounts are in $

Total assets carrying amount = 142,000 + 541,000 + 820,000 + 86,000 = 1,589,000

Total tax base = 108,000 + 340,000 + 610,000 + 40,000 = 1,098,000

Difference = 1,589,000 - 1,098,000 = 491,000

As tax base is lower than the carrying amount of assets, this results in creation of Deferred tax asset or reversal of Deferred tax liability.

Tax effect on such difference = 491,000 x 30%

= 147,300

Already existing Deferred tax liability is $170,000

Current year Tax payable is $8,400,000

Journal Entry

On June 2013

Income tax expense $8,252,700

Deferred tax liability $147,300

Incom Tax Payable $8,400,000

(Income tax payable for current year accounted along with reversal of Deferred tax liability arising due to temporary difference)

(b)

If there is existing Deferred tax liability is $120,000

Current year deferred tax asset or reversal of Deferred tax liability is $147,300

Of this we will use first $120,000 to reverse existing deferred tax liability and create new deferred tax asset for the remaining amount of $27,300

Journal entry

Income Tax Expense $8,252,700

Deferred tax liability $120,000

Deferred tax asset $27,300

Income tax payable $8,400,000

(Income tax liability for current year accounted along with the deferred tax asset arises due to temporary differences)